# Designing and Evaluating Hint Generation Systems for Science Education

### v1

@@ 3.2. Dataset
To evaluate the ability of large language models to generate hints, we use the SciQ dataset (Welbl et al., 2017a) for our evaluation. SciQ dataset is a science exam question answering dataset comprising of 13,679 multiple choice questions spread across biology, chemistry, physics and geology; spanning across elementary level to college introductory level questions. Since having multiple choices for each question hinders the ability to test the quality of generated hints, we discard the distractor choices, and view the task as a free-form short answer question answering task. We use two different dataset splits for the automatic and human evaluation as follows (see Table 1 for detailed statistics) -
@@
  • Dataset for human evaluation As the original SciQ dataset comprised of multiple-choice question answers, we hand-pick 30 questions from the test split of SciQ dataset to ensure valid open-ended questions with unambiguous answer, while ensuring equal representation of all four subjects in the final quiz.
3.3. Baselines
To extensively evaluate the capabilities of large language models to generate static and dynamic hints, we explore a diverse range of open-source model families. We limit our exploration on smaller open-source language models for three reasons - 1) open-source language model weights are constant, unlike API-based models, 2) there are privacy concerns with using closed-source models, and 3) any real-world application for learning feedback will have computational constraints, and thus we do not explore large language models that require multiple GPUs to execute. More specifically, in this work, we explore the following language models for the scientific hint generation task - DeepSeek-R1 (1.5B, 7B, 8B, 14B, 32B) (Guo et al., 2025), Gemma3 (1B, 4B, 12B, 27B) (Team et al., 2025), Mistral-Small (24B) [^1] https://huggingface.co/mistralai/Mistral-Small-24B-Instruct-2501, Phi4 (14B) (Abdin et al., 2024), and Qwen3 (0.6B, 1.7B, 4B, 8B, 14B, 30B, 32B) (Yang et al., 2025). We use an in-context learning approach to generate the static and dynamic chain-of-hints (refer to Figures 10 and 11 for the static and dynamic hint generation prompts respectively).
@@ 3.4. Automatic Evaluation Metrics
Reliable automatic evaluation metrics are critical in advancing research on automatic hint generation, and intelligent tutoring systems more broadly (Mousavinasab et al., 2021; Jangra et al., 2024). These automatic evaluation metrics provide a fast feedback cycle in early-stage system development. At the same time, they serve as a safeguard against premature deployment of underdeveloped systems. Exposing students to low-quality hint generation systems may not only undermine the learning outcomes, but also jeopardize learners’ trust in educational technologies more broadly. Having robust automatic evaluation metrics also helps ensure comparability and reproducibility of research findings across different studies. To automatically evaluate the pedagogical quality of the chain-of-hints in ways that align with learner needs and goals, we discuss several automatic evaluation metrics, either by adapting metrics from prior works, or proposing new ones wherever necessary. We briefly describe the evaluation metrics here, and provide detailed definitions in Appendix A.1.
We investigate a suite of evaluation metrics spanning five dimensions: information gain, redundancy, consistency, readability, and leakage. To capture the degree to which hints help learners progress toward correct answers, we propose information gain ($InfoGain$), which leverages a language model to compare question-answering performance with and without hints. We implement two variants: $InfoGain_{mean}$, the average gain across all hints in a chain, and $InfoGain_{comb}$, the maximum gain achieved by the entire hint chain.
Drawing on variation theory (Ling Lo, 2012), which emphasizes the value of presenting learners with diverse perspectives, we assess redundancy by adapting self-referenced redundancy from summarization research (Chen et al., 2021a), with the goal of minimizing repetitive information. To ensure factual correctness, we introduce consistency, computed using AlignScore (Zha et al., 2023), a metric designed to evaluate factual alignment between generated text and the source material.
@@
Experimental Setup. We evaluate all baselines over 1,000 instances from the validation split of the SciQ question answering benchmark (Welbl et al., 2017b). We explore two different hint generation settings to quantify the hint generation capabilities of these baselines - static and dynamic. For the static setting, we prompt the LLM baseline to generate 4 hints given a question answer pair. For the dynamic setting, we used the three distractor choices from the SciQ multiple choice questions as the incorrect responses, and interleaved them within each hint to replicate a learner interaction for the in-context learning prompt. We provide the prompts for static and dynamic hint generation in Figures 10 and 11 respectively.
Figure 2 summarizes the performance of the best-performing baselines from each model family. Comparing static and dynamic hint generation, we find that dynamic hints generally exhibit higher redundancy and lower consistency scores. This reduced content diversity, combined with elevated leakage rates, results in lower aggregate scores for dynamic hint chains relative to static hints. Across both settings, Mistral-Small-24b achieves the highest overall performance. Given the large number of baselines and evaluation settings, we report detailed experimental results in Appendix A.2. There, we examine trends across metrics, contrasts between static and dynamic strategies, and the impact of model parameters on hint quality. Later, in Section 5.4, we analyze the effectiveness of these automatic evaluation metrics to evaluate the quality of hints grounded on learner feedback from our user study.
@@ 4. User Study Design
We conducted a within-subjects human evaluation study with 41 participants to identify the characteristics of high-quality hints that learners might find most useful (RQ#1). We asked the participants to engage with our user interface in an unmoderated setting at their own convenience to obtain feedback in a more realistic self-study environment preventing any researcher bias.
@@
In the quiz assessment phase, the participants where given 30 free-form short answer questions standardized for all participants (8 physics, 8 biology, 7 chemistry and 7 geology questions). The quiz consisted of three sections, each containing 10 questions. With a within-subjects study design in mind, the first section served as the control condition with no hints provided, while the 2nd and 3rd sections featured either static or dynamic hints, assigned randomly to prevent order bias or recency bias in post-quiz briefing survey. For each question, participants had the option to request up to four hints before submitting an answer, and were allowed five attempts before their response was marked incorrect (see Figure 3). After each question, participants were presented with the correct answer and asked to evaluate the hints they had received. To assess the quality of each hint, we focused on three dimensions: participants’ overall satisfaction (five-point Likert scale), whether the hint provided new knowledge (informativeness), and whether it directly disclosed the correct answer (leakage). This structured feedback process allowed us to systematically assess participants’ perceptions of hint utility and informativeness, and assess the quality of our automatic evaluation metrics. After each section, participants were asked to answer four questions: one assessing the difficulty of the quiz on a five-point Likert scale, one rating the overall quality of hints in previous section on a five-point Likert scale, and two open-ended questions about the positives and negatives of hints shown in previous section. Due to lack of hints in the first Section, we only asked the first question in the 1st section’s survey. Participants progressed through the quiz at their own pace and had the option to take a break after completing each section, allowing them to pause and resume the study as needed.
@@
Participants. We recruited 46 participants by circulating our study through university mailing lists and slack channels across several universities in North America. We solicited participants who had studied science subjects in high school, lived in the United States, and were at least 18 years old. Each participant received a 20 dollars gift card for their participation. We do not include in our analysis the data from five participants that asked less than five hints to prevent introducing noise from learners who did not meaningfully engage with the system. From the remaining 41 participants in our study, we obtained feedback over 1,149 hints asked (639 static, 510 dynamic). 20 participants were first shown static hints and then dynamic hints, while for the other 21 participants the order was reversed to prevent ordering bias. On average, participants took 28.44 minutes to complete the study, where they asked 28.02 hints and successfully answered 21.85/30 questions. We provide the distribution of these statistics in Figure 4, and the aggregate participant demographics in Table 2. We redact all identifiable details from the data, and use participant IDs when quoting them in our findings.
@@ 5. Findings
In this section, we present findings from both qualitative and quantitative analyses. We begin with RQ#1, examining what makes hints useful to learners. This includes both the value of a single hint (Section 5.1) and the added complexity of chains of hints (Section 5.2). For RQ#2, we compare how learners perceive different hinting strategies, contrasting static and dynamic approaches (Section 5.3). Finally, for RQ#3, we address the alignment between automatic evaluation metrics and learners’ own judgments of hint quality (Section 5.4).
@@ 5.1. When a Single Hint Helps (or Hurts)
Our analysis reveals that at a sentence-level, hints portrayed several key characteristics of a successful feedback such as clarity and simplicity, yet recurring issues emerged around answer leakage and ambiguity. While most characteristics were clearly divided into good and bad across participants, we found that the preference was divided among the participants for certain types of hints. Our inductive analysis of learners’ feedback highlights several themes around perceived hint quality:
Improved conceptual understanding. Overall, participants marked 26.63% (306/1149) hints as informative, with a moderate correlation between the self-reported satisfaction and informativeness ($corr_{pearson}=0.4955$), and several participants explicitly noted that hints helped improve their understanding of the question, not just leading them to the answer. As P-22 described, ”hints encouraged to make associations and engage in reflection from a scientific perspective”. From out inductive analysis of qualitative responses, we identified three distinct ways in which hints supported this process. New perspectives: some hints reframed the questions to open new lines of thoughts. Linking to prior knowledge: other hints helped connect the new concepts to their own knowledge, helping them form associations. Supplementary knowledge: finally, several hints helped enrich participants’ knowledge with additional facts. Figure 5 (top left example) illustrates an interaction where hints help improve the participant’s understanding about the question by indicating a geological activity that involves magma rising to the surface to lead them to the correct answer volcanoes.
@@
In contrast, several participants emphasized that hints were unhelpful when they already recognized the concept but could not recall the precise term. As P-31 explained, ”If i remember the answer but it was on the tip of my tongue and cant remember the exact one, the hints do not help at all.”. This highlights a broader constraint of our short-form factoid question-answering setup: the system requires participants to supply the exact target term, meaning that hint effectiveness was often judged not only by whether it supported conceptual understanding, but also by whether it guided the learner toward the expected answer [^2] While we adopted a fuzzy match for our underlying answer assessment implementation, using in-context learning with an LLM (we used OpenAI’s gpt-3.5-turbo for the answer assessment task.), there were instances where it expected the exact lexical form (e.g., for the correct answer hydroxide ions, the submission OH- was marked incorrect). To ensure the efficacy of our assessment prompt for a functioning study, we test out the quality of the assessment model from a previous pilot study with 697 answers over the same 30 questions. While assessment model (see Figure 13 for the prompt) is able to achieve 88.95% accuracy over manually labeled responses, we delineate this limitation for real-world deployment in Section 6.1.1..
The tradeoff between goal reaching and knowledge gaining. Participants diverged in their views on answer-oriented hints helped them answer the question without improving conceptual understanding. Some appreciated strategies that narrowed the response space or partially disclosed the answer (e.g., the hint, ”The name starts with the letter A”, from Figure 5 bottom right interaction), even when these cues offered little additional conceptual knowledge. Others, however, criticized these hints as “misleading” or misaligned with their learning goals. As P-17 reflected, ”[…] A hint which helped me was good but it was not related to the concept itself. I prefer hints which use my external knowledge, but are also somewhat related to the question asked.” This divergent preference reflects a difference in underlying learning objectives (Hoque, 2016; Sönmez, 2017). For participants primarily aiming to maximize quiz performance, hints that directly enabled correct responses were welcomed. For others seeking to deepen their conceptual understanding, context-grounded hints were preferred—even at the cost of leaving the immediate question unanswered.
@@ 5.2. When One Hint Is Not Enough: The Dynamics of Chain-of-hints
While individual hints often provided clear and concise guidance, participants’ responses indicated that a single hint was not always sufficient to successfully answer the question. In many cases, participants required additional hints to fully grasp the concept, resolve ambiguities, or correct misconceptions. Across the 468 questions where hints were requested, only 32.5% were resolved with a single hint, whereas 21.5% required two hints, 13.9% required three, and 32.1% required four. These patterns suggest that learners frequently relied on sequences of hints to fully grasp the concept, resolve ambiguities, or correct misconceptions. Our analysis shows how the chain-of-hints shaped their overall effectiveness, revealing the conditions under which learners benefited from successive guidance as well as the pitfalls that emerged along the way.
@@
Figure 6. Distribution of hint chains (top) and corresponding satisfaction ratings for static and dynamic settings across attempted answers.

### v2

@@ 3.2. Dataset
To evaluate the ability of large language models to generate hints, we use the SciQ dataset (Welbl et al., 2017b) for our evaluation. SciQ dataset is a science exam question answering dataset comprising of 13,679 multiple choice questions spread across biology, chemistry, physics and geology; spanning across elementary level to college introductory level questions. Since having multiple choices for each question hinders the ability to test the quality of generated hints, we discard the distractor choices, and view the task as a free-form short answer question answering task. We use two different dataset splits for the automatic and human evaluation as follows (see Table 1 for detailed statistics) -
@@
  • Dataset for human evaluation As the original SciQ dataset comprised of multiple-choice question answers, we hand-pick 30 questions from the test split of SciQ dataset to ensure valid open-ended questions with unambiguous answers, while ensuring equal representation of all four subjects in the final quiz.
3.3. Models
To extensively evaluate the capabilities of large language models to generate static and dynamic hints, we explore a diverse range of open-source model families. We limit our exploration on smaller open-source language models for three reasons - 1) open-source language model weights are constant, unlike API-based models, 2) there are privacy concerns with using closed-source models, and 3) any real-world application for learning feedback will have computational constraints, and thus we do not explore large language models that require multiple GPUs to execute. More specifically, in this work, we explore the following language models for the scientific hint generation task - DeepSeek-R1 (1.5B, 7B, 8B, 14B, 32B) (Guo et al., 2025), Gemma3 (1B, 4B, 12B, 27B) (Team et al., 2025), Mistral-Small (24B) [^2] https://huggingface.co/mistralai/Mistral-Small-24B-Instruct-2501, Phi4 (14B) (Abdin et al., 2024), and Qwen3 (0.6B, 1.7B, 4B, 8B, 14B, 30B, 32B) (Yang et al., 2025). We selected these model families to capture a diverse range of open-source, instruction-tuned LLMs that vary in scale, training philosophy, and reasoning capability, while remaining practical for controlled experimentation. In particular, DeepSeek-R1 and Qwen3 include variants explicitly optimized for reasoning-style behavior, making them well-suited for evaluating multi-step hint generation. Gemma3, Mistral-Small, and Phi4 serve as strong general-purpose models with competitive performance across instruction-following and educational tasks. Together, these models provide broad coverage across parameter sizes (0.6B–32B) and architectural/modeling choices, enabling us to systematically study how model scale and reasoning specialization affect the quality and pedagogical structure of generated hint chains. We use an in-context learning approach to generate the static and dynamic chain-of-hints (refer to Figures 11 and 12 in Appendix B for the static and dynamic hint generation prompts respectively).
@@ 3.4. Automatic Evaluation Metrics
Reliable automatic evaluation metrics are critical in advancing research on automatic hint generation, and intelligent tutoring systems more broadly (Mousavinasab et al., 2021; Jangra et al., 2024). These automatic evaluation metrics provide a fast feedback cycle in early-stage system development. At the same time, they serve as a safeguard against premature deployment of underdeveloped systems. Exposing users to low-quality hint generation systems may not only undermine the learning outcomes, but also jeopardize their trust in educational technologies more broadly. Having robust automatic evaluation metrics also helps ensure comparability and reproducibility of research findings across different studies. To automatically evaluate the pedagogical quality of the chain-of-hints in ways that align with user’s needs and goals, we discuss several automatic evaluation metrics, either by adapting metrics from prior works, or proposing new ones wherever necessary. We briefly describe the evaluation metrics here, and provide detailed definitions in Appendix A.1.
We investigate a suite of evaluation metrics spanning five dimensions: information gain, redundancy, consistency, readability, and leakage. To capture the degree to which hints help users progress toward correct answers, we propose information gain ($InfoGain$), which leverages a language model to compare question-answering performance with and without hints. We implement two variants: $InfoGain_{mean}$, the average gain across all hints in a chain, and $InfoGain_{comb}$, the maximum gain achieved by the entire hint chain.
Drawing on variation theory (Ling Lo, 2012), which emphasizes the value of presenting learners with diverse perspectives, we assess redundancy by adapting self-referenced redundancy from summarization research (Chen et al., 2021b), with the goal of minimizing repetitive information. To ensure factual correctness, we introduce consistency, computed using AlignScore (Zha et al., 2023), a metric designed to evaluate factual alignment between generated text and the source material.
@@
Figure 2. Performance of best models from five baseline model families evaluated using automatic evaluation metrics. Orange (solid) bars denote the static hint generation results and blue (striped) bars denote the dynamic hint generation results. All metric values lie between 0 and 1, except for $Readability_{DC}$. Tables 4 and 4 provide detailed automatic evaluation metric results for the static and dynamic hint generation settings respectively.
3.5. Experimental Setup
We evaluate all baselines over 1,000 instances from the validation split of the SciQ question answering benchmark (Welbl et al., 2017a). We explore two different hint generation settings to quantify the hint generation capabilities of these baselines - static and dynamic. For the static setting, we prompt the LLM baseline to generate 4 hints given a question answer pair. For the dynamic setting, due to unavailability of user interaction data on the validation question-answer pairs, we used the three distractor choices from the SciQ multiple choice questions as the incorrect attempts, and interleaved them within each hint to replicate a learner interaction for the in-context learning prompt. For our human evaluation, however, we use actual responses to generate dynamic hints on-the-fly for better representation of a real-world interaction. We provide the prompts for static and dynamic hint generation in Appendix B, Figures 11 and 12 respectively.
3.6. Automatic Evaluation Results
Figure 2 summarizes the performance of the best-performing baselines from each model family. Comparing static and dynamic hint generation, we find that dynamic hints generally exhibit higher redundancy and lower consistency scores. This reduced content diversity, combined with elevated leakage rates, results in lower aggregate scores for dynamic hint chains relative to static hints. Across both settings, Mistral-Small-24b achieves the highest overall performance. Given the large number of baselines and evaluation settings, we report detailed experimental results in Appendix A.2. There, we examine trends across metrics, contrasts between static and dynamic strategies, and the impact of model parameters on hint quality. Later, in Section 5.4, we analyze the effectiveness of these automatic evaluation metrics to evaluate the quality of hints grounded on user feedback from our study.
@@ 4. User Study Design
We conducted a within-subjects human evaluation study with 41 participants to identify the characteristics of high-quality hints that learners might find most useful. We asked the participants to engage with our user interface in an unmoderated setting at their own convenience to obtain feedback in a more realistic self-study environment preventing any researcher bias.
@@
In the quiz assessment phase, the participants where given 30 free-form short answer questions standardized for all participants (8 physics, 8 biology, 7 chemistry and 7 geology questions). The quiz consisted of three sections, each containing 10 questions. With a within-subjects study design in mind, the first section served as the control condition with no hints provided, while the 2nd and 3rd sections featured either static or dynamic hints, assigned randomly to prevent order bias or recency bias in post-quiz briefing survey. At any point during a question, participants could request a hint or attempt an answer, with up to four hints and five attempts before their response was marked incorrect (see Figure 3). This open design allowed us to observe naturalistic help-seeking behavior and prevented the interface from artificially enforcing a fixed hint–attempt sequence.
After each question, participants were presented with the correct answer and asked to evaluate the hints they had received. To assess the quality of each hint, we focused on three dimensions: participants’ overall satisfaction (five-point Likert scale), whether the hint provided new knowledge (informativeness), and whether it directly disclosed the correct answer (leakage). We selected these three dimensions based on established principles in learning science and instructional design. Satisfaction reflects learners’ affective responses to support, which influence engagement (Ryan and Deci, 2000) and effective help-seeking (Aleven et al., 2004). Informativeness corresponds to whether the hint provides actionable scaffolding aligned with theories of feedback (Hattie and Timperley, 2007), cognitive load (Sweller, 1988), and the zone of proximal development (Vygotsky and Cole, 1978). Leakage captures the extent to which a hint over-assists by revealing the solution itself, directly relating to the assistance dilemma (Koedinger and Aleven, 2007) and research showing that excessive guidance undermines productive struggle (Renkl et al., 2004; Kapur, 2008). Together, these three dimensions map onto motivational, cognitive, and instructional considerations central to effective learning support, providing a theoretically grounded basis for evaluating hint quality.
After each section, participants were asked to answer four questions: one assessing the difficulty of the quiz on a five-point Likert scale, one rating the overall quality of hints in previous section on a five-point Likert scale, and two open-ended questions about the positives and negatives of hints shown in previous section. Due to lack of hints in the first Section, we only asked the first question in the 1st section’s survey. Participants progressed through the quiz at their own pace and had the option to take a break after completing each section, allowing them to pause and resume the study as needed.
@@
Participants. We recruited 46 participants by circulating our study through university mailing lists and slack channels across several universities. We solicited participants who had studied science subjects in high school, lived in the United States, and were at least 18 years old. Each participant received a 20 dollars gift card for their participation. We do not include in our analysis the data from five participants that asked less than five hints to prevent introducing noise from learners who did not meaningfully engage with the system. From the remaining 41 participants in our study, we obtained feedback over 1,149 hints asked (639 static, 510 dynamic). 20 participants were first shown static hints and then dynamic hints, while for the other 21 participants the order was reversed to prevent ordering bias. On average, participants took 28.44 minutes to complete the study, where they asked 28.02 hints and successfully answered 21.85/30 questions. Average self-reported quiz difficulty across all three sections on a 5-point Likert scale (1:very easy and 5:very hard) was 3.47/5, with a standard deviation of 0.47. We provide the distribution of these statistics in Figure 4, and the aggregate participant demographics in Table 2. We redact all identifiable details from the data, and use participant IDs when quoting them in our findings.
@@ 5. Findings
In this section, we present findings from both qualitative and quantitative analyses. We begin with RQ#1, examining what makes LLM-generated hints useful to participants. This includes both the value of a single hint (Section 5.1) and the added complexity of chains of hints (Section 5.2). For RQ#2, we compare how learners perceive different hinting strategies, contrasting static and dynamic approaches (Section 5.3). Finally, for RQ#3, we address the alignment between automatic evaluation metrics and learners’ own judgments of hint quality (Section 5.4).
@@ 5.1. When a Single Hint Helps (or Hurts)
Our analysis reveals that at a sentence-level, hints portrayed several key characteristics of a successful feedback such as clarity and simplicity, yet recurring issues emerged around answer leakage and ambiguity. While most characteristics were clearly divided into good and bad across participants, we found that the preference was divided among the participants for certain types of hints. Our inductive analysis of participant feedback highlights several themes around perceived hint quality:
Improved conceptual understanding. Overall, participants marked 26.63% (306/1149) hints as informative, with a moderate correlation between the self-reported satisfaction and informativeness ($corr_{pearson}=0.4955$), and several participants explicitly noted that hints helped improve their understanding of the question, not just leading them to the answer. As P-22 described, ”hints encouraged to make associations and engage in reflection from a scientific perspective”. Across all interactions, 109 hints directly assisted participants from an incorrect attempt to a correct submission. 83.4% of these hints were given high satisfaction ratings (4 or 5), and 59% of these hints were reported to be informative. We identified three distinct ways in which hints supported this process. New perspectives: some hints reframed the questions to open new lines of thoughts. Linking to prior knowledge: other hints helped connect the new concepts to their own knowledge, helping them form associations. Supplementary knowledge: finally, several hints helped enrich participants’ knowledge with additional facts. Figure 5 (top left example) illustrates an interaction where hints help improve the participant’s understanding about the question by indicating a geological activity that involves magma rising to the surface to lead them to the correct answer volcanoes.
@@
In contrast, several participants emphasized that hints were unhelpful when they already recognized the concept but could not recall the precise term. As P-31 explained, ”If i remember the answer but it was on the tip of my tongue and cant remember the exact one, the hints do not help at all.”. This highlights a broader constraint of our short-form factoid question-answering setup: the system requires participants to supply the exact target term, meaning that hint effectiveness was often judged not only by whether it supported conceptual understanding, but also by whether it guided them to the expected answer [^3] While we adopted a fuzzy match for our underlying answer assessment implementation, using in-context learning with an LLM (we used OpenAI’s gpt-3.5-turbo for the answer assessment task.), there were instances where it expected the exact lexical form (e.g., for the correct answer hydroxide ions, the submission OH- was marked incorrect). To ensure the efficacy of our assessment prompt for a functioning study, we test out the quality of the assessment model from a previous pilot study with 697 answers over the same 30 questions. While assessment model (see Figure 14 for the prompt) is able to achieve 88.95% accuracy over manually labeled responses, we delineate this limitation for real-world deployment in Section 6.2.1..
The tradeoff between goal reaching and knowledge gaining. Participants diverged in their views on utility of answer-oriented hints helped that them answer the question without improving conceptual understanding. Some appreciated strategies that narrowed the response space or partially disclosed the answer (e.g., the hint, ”The name starts with the letter A”, from Figure 5 bottom right interaction), even when these cues offered little additional conceptual knowledge. Others, however, criticized these hints as “misleading” or misaligned with their goals. As P-17 reflected, ”[…] A hint which helped me was good but it was not related to the concept itself. I prefer hints which use my external knowledge, but are also somewhat related to the question asked.” This divergent preference reflects a difference in underlying learning objectives (Hoque, 2016; Sönmez, 2017). For participants primarily aiming to maximize quiz performance, hints that directly enabled correct responses were welcomed. For others seeking to deepen their conceptual understanding, context-grounded hints were preferred—even at the cost of leaving the immediate question unanswered.
@@ 5.2. When One Hint Is Not Enough: The Dynamics of Chain-of-hints
While individual hints often provided clear and concise guidance, participants’ interactions indicated that a single hint was not always sufficient to successfully answer the question. In many cases, participants required additional hints to fully grasp the concept, resolve ambiguities, or correct misconceptions. Across the 468 questions where hints were requested, only 32.5% were resolved with a single hint, whereas 21.5% required two hints, 13.9% required three, and 32.1% required four. These patterns suggest that learners frequently relied on multiple hints to fully grasp the concept, resolve ambiguities, or correct misconceptions. Our analysis shows how the chain-of-hints shaped their overall effectiveness, revealing the conditions under which the participants benefited from successive guidance as well as the pitfalls that emerged along the way.
@@
(e) Percentage of successful static hint chains grouped by number of attempted answers.
(f) Percentage of successful dynamic hint chains grouped by number of attempted answers.
Figure 6. Distribution of hint chains (top), corresponding satisfaction ratings (middle) and percentage of successful interactions (bottom) for static and dynamic settings across attempted answers. Both static and dynamic hints have similar distribution of attempted answers, however, satisfaction of dynamic hints for instances without interactions is significantly lower than that of static hints. But as the number of attempted answers increase, the satisfaction ratings become similar. After four attempts, the interactions with four attempted answers yielded the least success rate compared to other interactions. This figure only considers interactions where hints were asked.